OpenStudy (anonymous):

One mole of an ideal gas is held at standard conditions. At what Kelvin temperature would the pressure be doubled?

OpenStudy (chmvijay):

keeping the volume constant right

OpenStudy (chmvijay):

PV=nRT P/T = constant P1/T2 =P2/T1 T2=P1 T1 /P2 but P2 is equal to 2P1 in this case T2=P1T1/2P1=1/2 T1 therefore temperature should be half of the initial temperature

OpenStudy (anonymous):

ok got it Thank you